2 ,3 ,7 ,29,?,3729
Answer by greenestamps(13200)



There is NEVER any way to know FOR SURE what the next number is in ANY problem like this. ANY number for the missing number makes a valid sequence.

However, given a problem like this, there might be a pattern that is relatively easy to see; in that case, it is LIKELY that the pattern you see is the "right" one. So on any problem like this I recommend you spend only a minute at most looking for a pattern and, if you don't see one in that amount of time, move on to some more productive activity.

In this example, I see...

3/2 is a bit more than 1
7/3 is a bit more than 2
29/7 is a bit more than 4

Looking at those numbers 1, 2, 4, ... leads to a pattern that is probably the right one:
      2 * 1 + 1 = 3
      3 * 2 + 1 = 7
      7 * 4 + 1 = 29
     29 * 8 + 1 = 233
    233 *16 + 1 = 3729

So it is LIKELY that the missing number is 233.